If a random variable x has the probability distribution arrayccccccccc x & 0 & 1 & 2 & 3 & 4 & 5 & 6 & 7 \ P(x) & 0 & 2k & k & 3k & 2k2 & 2k & k2 + k & 7k2 array then P(3 < x 6) is equal to
- 0.34
- 0.64
- 0.22
- 0.33
Answer
(4) 0.33
Solution
P(x) = 1
0 + 2k + k + 3k + 2k2 + 2k + k2 + k + 7k2 = 1
10k2 + 9k - 1 = 0
(10k - 1)(k + 1) = 0
k = 110 = 0.1 (k = -1 is rejected, as a probability cannot be negative)
P(3 < x 6) = P(4) + P(5) + P(6)
= 2k2 + 2k + k2 + k = 3k2 + 3k
= 0.03 + 0.3 = 0.33
Answer: (4)